# U-MedSAM: Uncertainty-aware MedSAM for Medical Image Segmentation

### v1

@@ 1 Introduction
Medical image segmentation is vital in clinical practice, enabling precise medicine [22], assessing therapeutic outcomes, and disease diagnosis by delineating organ boundaries and pathological regions, enhancing anatomical understanding, and abnormality detection [6]. Early segmentation models for medical images are typically designed for specific tasks with limited data. They might not handle the complex patterns and minute variations in medical images, which are often critical for clinical diagnosis and health science study [14].
@@ 2.2 Proposed Method
$\mathcal{L}=\frac{1}{2\sigma_{1}^{2}}\mathcal{L}_{1}+\frac{1}{2\sigma_{2}^{2}}\mathcal{L}_{2}+\cdots+\frac{1}{2\sigma_{m}^{2}}\mathcal{L}_{m}+\cdots\log\sigma_{1}\sigma_{2}\cdots\sigma_{m},$ (1)
where $M$ is the number of individual loss components, $\mathcal{L}_{m}$ represents each individual loss component (such as DC, CE, MSE, and Shape-Distance loss), and $\sigma_{m}$ are learnable noise parameters. Larger noise levels indicate that the talk is not properly understood, necessitating greater learning weights. These learnable parameters are optimized during the training process to minimize the overall loss. The last term can constrain the noise to be increased too much. Specifically, we incorporate the following loss functions, see the survey [7] to find more losses.
@@ 3.1 Dataset and evaluation measures
We only used the challenge dataset for model development. The evaluation metrics consist of two accuracy measures: the Dice Similarity Coefficient (DSC) and Normalized Surface Dice (NSD), along with an efficiency measure of running time. Together, these metrics contribute to the overall ranking computation.
@@ 4.2 Qualitative results on validation set
We show some examples with good segmentation results in Fig. 3 and examples with bad segmentation results in Fig. 4.

### v2

@@ 1 Introduction
Medical image segmentation is vital in clinical practice, enabling precise medicine [24, 23], assessing therapeutic outcomes, and disease diagnosis by delineating organ boundaries and pathological regions, enhancing anatomical understanding, and abnormality detection [6]. Early segmentation models for medical images are typically designed for specific tasks with limited data. They might not handle the complex patterns and minute variations in medical images, which are often critical for clinical diagnosis and health science study [14].
@@ 2.2 Proposed Method
2.2.1 Knowledge Distillation
In the LiteMedSAM model, both the prompt encoder and mask decoder are lightweight models, and the image encoder compresses the model through TinyViT. However, TinyViT uses Softmax-based self-attention mechanism, and the computational complexity is still high. So, in order to reduce the computational complexity of the model, we use EfficientViT to replace TinyViT. EfficientViT reduces the computational complexity from quadratic to linear by lightweight ReLU linear attention. During the initialisation process, we use TinyViT as a teacher model for knowledge distillation, efficiently migrated to EfficientViT. Prompt encoder and mask decoder are already very lightweight models and no other changes were made. The knowledge distillation process was supervised using the L2 loss function.
@@ 2.2.1 Knowledge Distillation
$\mathcal{L}=\sum_{m=1}^{M}\left(\frac{1}{2\sigma_{m}^{2}}\mathcal{L}_{m}+\log(1+\sigma_{m}^{2})\right)$ (1)
where $M$ is the number of individual loss components, $\mathcal{L}_{m}$ represents each individual loss component (such as DC, CE, MSE, and Shape-Distance loss), and $\sigma_{m}$ are learnable noise parameters. Larger noise levels indicate that the talk is not properly understood, necessitating greater learning weights. These learnable parameters are optimized during the training process to minimize the overall loss. The last term can constrain the noise to be increased too much [20]. Specifically, we incorporate the following loss functions, see the survey [7] to find more losses.
@@ 3.1 Dataset and evaluation measures
We only used the challenge dataset for model development.
The evaluation metrics consist of two accuracy measures: the Dice Similarity Coefficient (DSC) and Normalized Surface Dice (NSD), along with an efficiency measure of running time. DSC is used to measure the degree of overlap between the segmentation result and the ground truth. And NSD is a surface-based distance metric for measuring the difference between segmentation boundaries, which takes into account the distance between the boundary of the segmentation result and the label boundary. They are calculated as follows in equations 2 and 3:
$\mathrm{DSC}=\frac{2\times|y\cap\hat{y}|}{|y|+|\hat{y}|}$ (2)
$\mathrm{NSD}=\frac{|\{x\in y:d(x,\hat{y})\leq T\}|+|\{x\in\hat{y}:d(x,y)\leq T\}|}{|y|+|\hat{y}|}$ (3)
where, $y$ and $\hat{y}$ represent the segmentation result and the ground truth, respectively. $d(x,A)$ denotes the shortest distance from point $x$ to the surface of set $A$, and $T$ is the tolerance distance threshold. Both DSC and NSD are closer to 1, indicating better segmentation performance, which implies a higher degree of overlap and boundary similarity between the predicted segmentation and the ground truth annotation. Together, these metrics contribute to the overall ranking computation.
@@ 4 Results and discussion
Figure 5: Scribble-based cases with good results for various modality segmentations.
@@ 4.2 Qualitative results on validation set
We show some examples with good segmentation results in Fig. 3 and examples with bad segmentation results in Fig. 4. Fig. 5 and Fig. 6 show the results of good and poor visualisation of scribble-based segmentation, respectively.
Figure 6: Scribble-based cases with bad results for various modality segmentations.